OpenStudy (anonymous):

Solve the initial value problem analytically: dy/dx = 1 + x + (x^2/2) y(0)=1

OpenStudy (perl):

did you try integrating

OpenStudy (anonymous):

I know you need to integrate but I think that's where I'm having problems. The answer in the back of the book is not at all like my answer.

OpenStudy (perl):

$$\large \frac{dy}{dx} = 1 + x + \frac{x^2}{2} ~~,~~y(0)=1\\ \large y(x) = x + \frac{x^2}{2} + \frac{1}{2}\cdot \frac{x^3}{3} + C \\ \large y(0) = 0 + \frac{0^2}{2} + \frac{1}{2}\cdot \frac{0^3}{3} + C \\ \large y(0) = 1 \\ \large \iff \\ C = 1 $$

OpenStudy (anonymous):

Oh so dx just means I need to find the antiderivative of that. Got it that's what was confusing me I guess. Okay thank you again Perl. I've got a test tomorrow and you may have saved my life

OpenStudy (perl):

your welcome